OpenStudy (anonymous):

The ratio of the lengths of the corresponding sides of two rectangles is 8:3. The area of the larger rectangle is 320 ft2. What is the area of the smaller rectangle?

OpenStudy (er.mohd.amir):

let side of 1 rectangle be 8x then second is 3x and other side is 8y and 3y larger area=8x*8y=320 64 *x*y=320 x*y=320/64 small area=3x*3y =9*x*y=9*320/64 ans

OpenStudy (er.mohd.amir):

45 feet2